# Read, Watch and Scream! Sound Generation from Text and Video

### v1

@@ 3.2 Rethinking video-to-sound generation with temporal alignment
We hypothesize a video input can bring principal information missing in a text prompt, i.e., temporal alignment for a visual scene. Prior sound generation methods from visual content were immersed in the overall audiovisual relevance [3, 34, 38]. Since then, recent works [7, 4, 43] have attempted to generate SFX for video by estimating the onset [4], or audio timestamp [43] from videos. However, these methods produce an unnatural sound in that discrete conditions such as onset or timestamp cannot serve continuous sound variations.
@@
In this work, we consider energy, interpreted as “loudness” or “dynamics”, to produce a continuous condition. Figure 3 shows that energy is a continuous time-varying signal, including envelope components of sound such as peak, attack, sustain, and decay. Energy can be obtained cheaply and automatically by computing the frame-level magnitude of mel-spectrograms [31]. Moreover, we empirically observe that energy can also implicitly imply multiple semantics of the video. For example, Figure 3 shows the generated audio by ReWaS contains multiple semantics beyond the text control “playing darts”; it generates people talking sounds for the intervals having less power than the sound of a dart hitting the target. In other words, the energy information can imply the existence of different objects in the video as well as the main object.
@@ 4 Method
In this work, we introduce a novel sound generation method conditioned on both text and video, to generate a waveform that is temporally well aligned with the visual input. As shown in Figure 4, our model consists of two main parts: (i) control prediction module, which intermediately predicts energy control from the video. (Section 4.1) (ii) conditional sound generation module, which uses the energy control signal as a condition in the diffusion process to generate corresponding audio outputs (Section 4.2), which are both temporally and semantically aligned with text and video.
@@
Figure 4: Overall architecture of ReWaS. Our model predicts energy control from a given video, and generates sound with text prompt and control condition. Red lines are used in training only.
4.1 Control prediction from video
@@
Energy control. ReWaS is based on a T2A generation method, specifically AudioLDM that uses CLAP embedding space for text and audio alignment. A naïve approach using video as a condition is to align latent space between audio-video-text. Luo et al. [27] attempted to align tri-modal embeddings in a unified space by large-scale contrastive learning prior to training diffusion models. To more efficiently overcome this challenge, we design an energy control as an intermediate bridge from video to audio. We speculate that energy control brings three advantages: First, the power of audio is intuitively correlated to visual dynamics and semantics [20, 37]. With the natural fact that people can imagine the power of sound from the size of the instance or distance to the object, we regard audio energy as a visually correlated signal that can be certainly obtained from video. Second, as shown in previous works [31, 13], energy plays as a structural condition for audio generation. Therefore, it is well-suited to parameter-efficient fine-tuning methods such as ControlNet [46]. Finally, using temporal acoustic signals for generating audio needs a skilled user to annotate the pitch, melody, or rhythm for every timestamp. It makes the audio generation phase impractical and difficult for the public to control. Meanwhile, energy is highly related to physical interactions implicated in visual signals; thus, it can be easily estimated from the video. Our approach does not require timestamp-wise fine-grained user control, but automatically estimating energy structure from the given video.
@@
Video embedding. To predict the energy control from video input, we use the feature extracted from a pretrained SynchFormer [19] video encoder. We empirically observe that the image encoder (e.g., CLIP [30]) is limited to V2A generation, especially from a temporal alignment perspective. We finally take video embedding $\mathbf{E}_{v}\in\mathbb{R}^{S\times C}$, where $S$ is the number of segments and $C$ is the dimension of latent. The implementation details for this process are described in Appendix A.1.
@@
Training energy control from video. Similar to Ren et al. [31], we calculate the energy from the mel-spectrogram by averaging the frequency bins and further smoothing the time-sequential energy information. We first transform the raw waveform to the mel-spectrogram, $\texttt{mel}\in\mathbb{R}^{D\times W}$, where $D$ represents the number of mel-frequency bins, and $W$ is the width of the spectrogram following AudioLDM [25]. However, we empirically observe that the computed energy fluctuates a lot for each temporal window, which hinders stable training. We resolve the issue by taking a smoothing operator. The energy of audio $e\in\mathbb{R}^{W}$ is defined as $e_{a}=\texttt{Smoothing}\left(\frac{1}{D}\sum_{d=1}^{D}\texttt{mel}_{w,d}\right)$. We use the second-order Savitzky-Golay filter [40] with a window length of 9 for smoothing.
@@ 4.2 Conditional sound generation
Adding control signal to diffusion model. To reflect the energy control signal, we train the energy adapter following the framework of ControlNet [46]. The weights of the energy adapter are initialized from pretrained parameters of diffusion models, and connected to AudioLDM with zero convolution layers. Compared to training audiovisual alignment into the latent space in diffusion model [27, 44], our adapter takes the benefit of robust fine-tuning speed (e.g., [27] uses 8 A100 GPUs for 140 hours for feature alignment and LDM tuning, whereas ReWaS uses 4 V100 GPUs for total 33 hours). To add the control feature for $z_{t}$, the energy control $e_{a}$ is duplicated by the number of mel-filterbanks, and transferred to the VAE encoder for the purpose of encoding, followed by a fully-connected layer and SiLU activation [8]. This latent control feature $c_{e}$ is added to the $z_{0}$, where $z_{0}$ is an audio prior obtained from the VAE encoder. Thus, given a text embedding $\mathbf{E}_{y}$ and latent control feature $c_{e}$, we train energy adapter by optimizing the following objective: $\mathcal{L}_{c}=\mathbb{E}_{z_{0},t,\mathbf{E}_{y},c_{e},\epsilon\sim\mathcal{N}(0,1)}\|\epsilon-\epsilon_{\theta}(z_{t},t,\mathbf{E}_{y},c_{e})\|_{2}^{2}.$ During training, we randomly drop $\mathbf{E}_{y}$ with the probability 0.3 for better controls. We denote that $\mathcal{L}_{c}$ and $\mathcal{L}_{e}$ are optimized separately.
Sound generation. We employ DDIM sampler [36] to generate sound from the noise. The reverse sampling process is conditioned on both text prompt and video. We replace $e$ to $\hat{e}=\phi(\mathbf{E}_{v})$ at inference. Once the mel spectrogram is generated by the VAE decoder, it can be transformed into a raw waveform using a pretrained neural vocoder [23] as explained in Section 3.1.
@@ 5.1 Experimental settings
Datasets. To ensure a fair comparison with existing baselines, we train both the control prediction module and the adapter in the conditional sound generation module on VGGSound [2]. VGGSound is a large-scale audiovisual dataset containing approximately 200k video clips of 10-second duration, accompanied by corresponding audio tracks. The dataset covers 309 classes of general sounds, roughly categorizing them into acoustic events, music, and people. The videos are sourced from YouTube, providing a diverse and realistic corpus. For our experiments, we leverage a subset of 160k videos from VGGSound due to the availability of public videos at the time of training. Since the VGGSound includes plentiful general sound examples, our method trained on the VGGSound enables general-purpose sound generation for real-world scenarios. We split the train data list into training and validation subsets following SpecVQGAN [18]. To evaluate temporal alignment accuracy, we evaluate ReWaS on Greatest Hits [28] test set including the videos of hitting a drumstick with materials. Since Greatest Hits samples have a distinct audio property compared to the other audio samples, we fine-tune our method on the Greatest Hits training samples. To the best of our knowledge, it is the first proposed evaluation protocol to evaluate the model on both of these datasets.
Evaluation protocol. In all experiments, we generate 5 seconds of audio corresponding to the input video clips and text prompt. We use sound class categories as the text prompt. We observe that the existing baselines used different numbers of generated samples and audio sequence time, which could lead to inconsistencies in the evaluation. For a fair comparison, we re-implement all baseline methods to generate one audio sample with 5 seconds from one video in VGGSound dataset, consistent with ReWaS. On the other hand, 2 seconds of generation results are used in Greatest Hits.
@@
Baselines. We compare ReWaS against existing video-to-sound generation approaches in priority, SpecVQGAN [18], Im2wav [34], and Diff-Foley [27], which are trained on the VGGSound and AudioSet datasets. For a fair comparison, we take the following steps: We first generate the full-length audio by SpecVQGAN and Diff-Foley (10s and 8s, respectively) for the test videos. Then, we extract the 5-second clip corresponding to the same video frames used in our method. Since Im2wav is designed to generate sound with a fixed length of 4 seconds, we first generate the initial 4 seconds and extend it by generating an additional 1 second, resulting in a 5-second audio clip. In the temporal alignment evaluation, we consider CondFoleyGen [7] as a main baseline, which is trained on the Greatest Hits dataset.
@@
Evaluation metrics. To evaluate perceptual quality and diversity, we employ Fréchet distance (FID) [14] and inception score (IS) [33]. IS assesses sample quality and diversity, while FID measures distribution-level similarity. We also report the mean of KL divergence (MKL) [18] for paired samples to evaluate the relevance to the condition. Following the implementation of SpecVQGAN [18], we use a Melception classifier to report those scores. We also measure the alignment between the generated audio and sound categories with CLAP score [17] in VGGSound. In the Greatest Hits experiment, we report onset accuracy (Acc) and average precision (AP), following the evaluation protocol introduced by CondFoleyGen [7]. The onset of sound events is a discrete signal obtained by the thresholding of the amplitude gradient. Therefore, it is challenging to detect onsets in natural sound like VGGSound benchmark. To address this issue, we report the mean absolute error (MAE), following the approach in  [12], to evaluate the difference energy signal from real and generated sounds for the first time in the sound generation task conditioned on video. Although these evaluation metrics can evaluate different properties of the generated audio, all of them measure the difference between the generated audio and the “ground truth” audio corresponding to the original video. However, a sound is multimodal information; one video can sound differently (e.g., the voice of a human can vary). Consequently, existing quantitative evaluation metrics have challenges in measuring whether the generated audio is truly suited to the given video. To tackle the issue, we conduct a user study to evaluate the quality of the generated audio samples. More details are explained in the next subsection.
@@ 5.2 Results
Table 2: Performance comparison on VGGSound [2] with reproduced five seconds audio samples.
@@
SpecVQGAN [18] 38.88 7.41 7.22 12.34 2.27 379M
Im2wav [34] 33.10 10.72 5.90 25.70 3.31 365M
Diff-Foley [27] 30.53 17.03 6.61 38.11 2.08 859M
ReWaS (Ours) 29.09 15.96 6.37 41.70 1.29 204M
Table 3: Performance comparison for sound generation on Greatest Hits [28]. All baselines are evaluated on re-produced results. denotes the models trained on Greatest Hits. We use material types as text prompts, while CondFoleyGen uses both reference audio and video as inputs.
@@
SpecVQGAN [18] 42.72 1.57 6.13 18.09 62.78 4.29
Im2wav [34] 85.64 1.64 4.83 14.36 60.44 3.77
Diff-Foley [27] 86.40 1.74 5.07 23.94 65.61 3.40
CondFoleyGen [7] 40.89 1.63 7.08 23.94 60.24 2.94
ReWaS (Onset) 39.96 3.09 7.02 21.81 65.88 3.64
ReWaS (Energy) 36.70 4.85 3.92 19.15 63.28 2.93
Quantitative results. Table 2 shows the quantitative comparisons on the VGGSound benchmark. We note that category classes are used as text prompts in the VGGSound. Compared to GAN-based SpecVQGAN and language model-based Im2wav, the models based on diffusion (Diff-Foley and ReWaS) show high-fidelity sounds (i.e., better FID, IS and MKL). We train 22M parameters for video projection to audio conditional control, and 182M parameters for fine-tuning the AudioLDM with our energy adapter. Especially, while we use only a quarter of training parameters compared to Diff-Foley [27], our method outperforms Diff-Foley on all metrics except IS score. CLAP scores illustrate the importance of text prompts for semantic alignment. Also, our method surprisingly outperforms in terms of MAE. This result demonstrates the accuracy of our control prediction module, and generated outputs from ReWaS are most temporally closer to the real content.
In addition, we evaluate how the generated audio and the condition video are temporally aligned on Greatest Hits. The dataset distribution of Greatest Hits highly differs from the general audio samples; hence, we fine-tune ReWaS on the Greatest Hits training samples. Table 3 shows the results. ReWaS achieves the best audio generation quality (FID, IS, MKL), and reasonable onset accuracy and AP, although ReWaS is not specially designed for Foley like Diff-Foley and CondFoleyGen. Finally, our method shows the best energy prediction score which forms the overall sound structure.
We also implement ReWaS with the onset signal, which is a discrete sequence to represent 0 or 1 (details can be found in Section A.1 and A.2). Compared to ReWaS with energy, our onset model shows more accurate results for temporal alignment metrics. It demonstrates that although the onset signal is advantageous in predicting the exact starting point of the sound, the energy is advantageous in terms of the overall audio quality and the homogeneity of the sound (i.e., energy MAE).
Table 4: Human evaluation of V2A methods for comparison of audio quality, audiovisual relevance, and temporal alignment with 5-scale MOS.
@@
SpecVQGAN [18] 2.76 2.50 2.64
Im2wav [34] 2.97 3.18 3.01
Diff-Foley [27] 2.89 2.97 2.98
@@
User study. The quantitative results are limited to measuring how the generated audio sounds realistic and aligned to the given video. To complement the quantitative evaluations, we conduct a human evaluation study to assess the subjective quality of the generated audio concerning the input video. We ask the human evaluators to evaluate the quality of the audio samples generated by SpecVQGAN [18], Im2wav [34], Diff-Foley [27], and ReWaS.
We use three evaluation criteria: audio quality, relevance between audio and video, and temporal alignment. Detailed user instructions are depicted in Section A.3. We use a five-point Likert scale to measure mean opinion score (MOS), where an ideal video clip with its ideal audio receives a rating of 5 across all criteria. We recruit human annotators via two separate channels: Amazon Mechanical Turk (AMT) and local hiring. For AMT, we recruit 50 annotators for each criterion, and each annotator evaluates five generated samples for each method (i.e., each annotator evaluates 20 generated audios). For locally hired annotators, we ask them to evaluate 20 generated samples for each method and criterion. There were 23 local annotators, and each of them evaluated 240 samples.
Table 4 shows that ReWaS achieves the best in all categories. This subjective human evaluation result is consistent with our quantitative and qualitative findings, further validating the effectiveness of our approach in generating high-quality, relevant, and temporally synchronized audio for the given video.
Figure 5: ReWaS successfully generates corresponding audio compared to reference (ground-truth) audio. Surprisingly, when the skateboarder jumps into the air, only ReWaS succeded in detecting short transition moments (yellow box). The input text prompt is “skateboarding”.
Qualitative results. Figure 5 shows qualitative results in baselines and ReWaS. Given the skateboarding video, SpecVQGAN and Diff-Foley fail to generate the sound of skate wheels rolling on the floor. Although Im2wav generates that sound, it cannot capture a short transition.
We also demonstrate the effectiveness of the text prompt in Figure 6 with CLAP similarity. In the input video shown in Fig. 6 (a), rain streaks are barely visible, while we want to hear the sound of rain. ReWaS can emphasize the desired sound with the help of a text prompt ‘raining’. Meanwhile, videos in the wild may include redundant frames as shown in Fig. 6 (b). In this case, V2A methods also struggle to generate corresponding sound. However, ReWaS can effectively calibrate the semantics by user prompt. More examples can be found on our demo page.
Table 5: Impact of the energy control’s quality. Numbers are measured on VGGSound.
@@
Text & the ground-truth audio energy (upper bound) 29.06 18.21 6.12 44.97 0.95
Text & the estimated energy from the video 29.09 15.96 6.37 41.70 1.29
5.3 Discussion and limitation
The impact of the quality of the energy control. ReWaS estimates the audio energy from the video. However, the estimated energy would propagate an estimation error, which may cause a significant performance drop. To verify the robustness of the energy prediction module, we compare the control by our video-to-energy prediction module and the energy directly extracted from the ground truth audio. Table 5 shows that although we use the estimated energy, the quality of the generated audio is very similar to the audio samples controlled by the ground truth audio energy. It supports the idea that energy information is highly related to visual information and is easy to estimate solely using video.
Figure 6: Effectiveness of text prompt. Videos in the real world are sometimes noisy. For example, when videos (a) are hard to distinguish the semantics or (b) contain redundant frames, text prompts used in ReWaS calibrate the results. Text prompt in (a) is “raining”, and (b) is “chicken clucking”.
Figure 7: Effectiveness of video input. In ReWaS, multiple semantics in the video are transferred through energy control. Even if there is missing information in the text prompt, the energy control complements this.
Effectiveness of video input. Figure 7 illustrates examples including multiple semantics. For instance, when a person talking and playing a dart game in an input video, the original AudioLDM generates only the sound of talking, ignoring ‘dart’ prompt. Additionally, aligning generated sound with video is challenging in AudioLDM. In comparison, ReWaS not only generates both the sound of talking and dart but also aligns the sound with the frames. Since the energy encompasses not just volumes but also the power of acoustical features (frequency), this supports the notion that multiple semantics can also be conveyed by our energy condition. Furthermore, the result demonstrates the limitation of T2A methods for automatic Foley synthesis, because they cannot watch a video. Another example is in Figure A.5.
Limitation. Although our approach successfully leverages the text and video control simultaneously, our method shares the limitation of AudioLDM, namely, hallucination in generated samples. For example, for a given “basketball bounce” video, ReWaS generates a squeaking sound, even if the player is standing still. This problem might be mitigated if we can use a better AudioLDM model.

### v2

@@ 3.2 Video-to-audio with temporal alignment
We assert that a video input can bring principal temporal information that is hard to convey with a text prompt. However, directly injecting temporal information from visual into an audio generation model remains a significant challenge. In contrast, previous works have attempted to generate sound by estimating the onset (Comunità et al. 2024), or audio timestamp (Xie et al. 2024) from videos to improve audiovisual relevance. However, they are limited to producing an unnatural sound for a single object in that discrete conditions cannot serve continuous sound variations.
@@
In this work, we consider energy, the averaged mel-spectrogram on the frequency axis, to produce a continuous condition. Figure 3 shows that energy is a continuous time-varying signal, including envelope components of sound such as peak, attack, sustain, and decay. Energy can be obtained cheaply and automatically by computing the frame-level magnitude of mel-spectrograms (Ren et al. 2020). Moreover, we empirically observe that energy can also implicitly improve the temporal alignment of the video. For example, Figure 3 shows energy can contain continuously varying audio information.
@@ 4 Method
This paper introduces a novel sound generation method conditioned on text and video, to generate a waveform temporally well aligned with the visual input. Our model consists of two parts: (i) control prediction, which intermediately predicts energy control from the video. (Section 4.1) (ii) conditional sound generation, which uses the energy control signal as a condition in the diffusion process to generate corresponding audio outputs (Section 4.2), which are both temporally and semantically aligned with text and video.
@@
Figure 4: Overall architecture of ReWaS. Our model predicts energy control from a given video, and generates sound with text prompt and control condition. Red lines are used in training only, and replaced to the video-to-energy estimator $\phi$ in test time.
4.1 Energy control prediction from video
@@
Energy control. ReWaS is based on AudioLDM that uses CLAP embeddings for text and audio alignment. A naïve approach using video as a condition is to align latent space between audio-video-text. Luo et al. (2024) attempted to align tri-modal embeddings in a unified space by large-scale contrastive learning prior to training diffusion models. To more efficiently overcome this challenge, we design an energy control as an intermediate bridge from video to audio. We speculate that energy control brings three advantages: First, the power of audio is intuitively correlated to visual dynamics and semantics (Jeong et al. 2023; Sung-Bin et al. 2023). With the natural fact that people can imagine the power of sound from the size of the instance or distance to the object, we regard audio energy as a visually correlated signal that can be certainly obtained from video. Second, as shown in Ren et al. (2020) and Guo et al. (2024b), energy plays as a structural condition for audio generation. Thus, it is well-suited to parameter-efficient fine-tuning methods such as ControlNet. Finally, using temporal acoustic signals for generating audio needs a skilled user to annotate the pitch, melody, or rhythm for every timestamp. It makes the audio generation phase impractical and difficult for the public to control. Meanwhile, energy is highly related to physical interactions implicated in visual signals; thus, it can be easily estimated from the video. Our approach does not require timestamp-wise fine-grained user control, but automatically estimating energy structure from the given video.
@@
Video embedding. To predict the energy control from video input, we extract features from the pretrained SynchFormer (Iashin et al. 2024) video encoder. We empirically observe that the image encoder (e.g., CLIP (Radford et al. 2021)) is limited to V2A generation, especially from a temporal alignment perspective. We finally take video embedding $\mathbf{E}_{v}\in\mathbb{R}^{S\times C}$, where $S$ is the number of segments and $C$ is the dimension of latent. The implementation details for this process are described in Appendix.
@@
Training energy control from video. Similar to Ren et al. (Ren et al. 2020), we calculate the energy from the mel-spectrogram by averaging the frequency bins and further smoothing the time-sequential energy information. We first transform the raw waveform to the mel-spectrogram, $\texttt{mel}\in\mathbb{R}^{D\times W}$, where $D$ represents the number of mel-frequency bins, and $W$ is the width of the spectrogram following AudioLDM (Liu et al. 2023). However, we empirically observe that the computed energy fluctuates a lot for each temporal window, which hinders stable training. We resolve the issue by taking a smoothing operator. The energy of audio $e\in\mathbb{R}^{W}$ is defined as $e_{a}=\texttt{Smoothing}\left(\frac{1}{D}\sum_{d=1}^{D}\texttt{mel}_{w,d}\right)$. We use the second-order Savitzky-Golay filter (Virtanen et al. 2020) with a window length of 9 for smoothing.
@@ 4.2 Conditional sound generation
Adding control signal. To reflect the energy control signal, we train the energy adapter following ControlNet (Zhang, Rao, and Agrawala 2023). The weights of the energy adapter are initialized from pretrained parameters of diffusion models, and connected to AudioLDM with zero convolution layers. Compared to training audiovisual alignment into the latent space in diffusion model (Luo et al. 2024; Xing et al. 2024), our adapter takes the benefit of robust fine-tuning speed (e.g., Luo et al. (2024) uses 8 A100 GPUs for 140 hours for feature alignment and LDM tuning, whereas we use 4 V100 GPUs for total 33 hours). To add the control feature for $z_{t}$, the energy control $e_{a}$ is duplicated by the number of mel-filterbanks, and transferred to the VAE encoder for the purpose of encoding, followed by a fully-connected layer and SiLU (Elfwing, Uchibe, and Doya 2018). This latent control feature $c_{e}$ is added to the $z_{0}$, where $z_{0}$ is an audio prior obtained from the VAE encoder. Thus, given a text embedding $\mathbf{E}_{y}$ and latent control feature $c_{e}$, we train energy adapter by optimizing the following objective: $\mathcal{L}_{c}=\mathbb{E}_{z_{0},t,\mathbf{E}_{y},c_{e},\epsilon\sim\mathcal{N}(0,1)}\|\epsilon-\epsilon_{\theta}(z_{t},t,\mathbf{E}_{y},c_{e})\|_{2}^{2}.$ During training, we randomly drop $\mathbf{E}_{y}$ with the probability 0.3 for better controls. We denote that $\mathcal{L}_{c}$ and $\mathcal{L}_{e}$ are optimized separately.
Sound generation. We use DDIM (Song, Meng, and Ermon 2020) to generate sound from the noise. The reverse sampling process is conditioned on both text and video. We replace $e$ to $\hat{e}=\phi(\mathbf{E}_{v})$ at inference. Once mel-spectrogram is generated by the VAE decoder, it can be transformed into a raw waveform using the pre-trained vocoder (Kong, Kim, and Bae 2020) (See Section 3.1). Note that conditioning on video increases the total inference time by only 3%.
@@ 5.1 Experimental settings
Datasets. For a fair comparison with existing baselines, we train the control prediction module and the adapter in the conditional sound generation module on VGGSound (Chen et al. 2020). VGGSound is a large-scale dataset containing $\approx$200k video clips, accompanied by corresponding audio tracks. The dataset covers 309 classes of general sounds, roughly categorizing them into acoustic events, music, and people. The videos are sourced from YouTube, providing a diverse and realistic corpus. Since the VGGSound includes plentiful general sound examples, ReWaS trained on the VGGSound enables general-purpose sound generation for real-world scenarios. We randomly sampled 3K videos to construct VGGSound test subset. To evaluate temporal alignment accuracy, we use Greatest Hits (Owens et al. 2016) test set including the videos of hitting a drumstick with materials. Since Greatest Hits samples have a distinct audio property compared to the other audio samples, we fine-tune ReWaS on the Greatest Hits training samples.
@@
Baselines. We compare ReWaS against open-source V2A generation approaches in priority, SpecVQGAN (Iashin and Rahtu 2021), Im2wav (Sheffer and Adi 2023) and Diff-Foley (Luo et al. 2024), which are trained on the VGGSound and AudioSet datasets. Furthermore, we compare Seeing&Hearing (Xing et al. 2024), which optimizes a pre-trained AudioLDM during the inference stages by aligning the latent space using ImageBind. For a fair comparison, we take the following steps: We first generate the full-length audio by each method, and use a common 5-second clip for evaluation. In the temporal alignment evaluation, we consider CondFoleyGen (Du et al. 2023) as a main baseline, which is trained on the Greatest Hits dataset.
@@
Evaluation metrics. Following the implementation of AudioLDM, we employ Fréchet distance (FD) (Heusel et al. 2017), Fréchet audio distance (FAD) (Tailleur et al. 2024), and the mean of KL divergence (MKL) (Iashin and Rahtu 2021). We also measure the alignment between the generated audio and sound categories with CLAP score (Huang et al. 2023b). However above metrics are limited to evaluating audio-visual temporal alignment, so we employ AV-align (Yariv et al. 2024) based on detecting energy peaks in audio-visual modalities. In the Greatest Hits experiment, we report onset accuracy (Acc) and average precision (AP), following the evaluation protocol introduced by CondFoleyGen. The onset of sound events is a discrete signal obtained by the thresholding of the amplitude gradient. Therefore, relatively quiet sound effects (e.g., scratching leather, touching the leaves) or natural sounds can be excluded from the evaluation. To address this issue, we report the mean absolute error (MAE) (Guo et al. 2024a) of the energy signals from real and generated sounds for the first time in the sound generation task conditioned on video. Although these evaluation metrics can evaluate different properties of the generated audio, most of them measure the difference between the generated audio and the “ground truth” audio corresponding to the original video. However, one video can sound differently (e.g., human’s voice can vary); existing quantitative evaluation metrics have challenges in measuring whether the generated audio is truly suited to the given video. To tackle the issue, we conduct a user study to evaluate the quality and temporal alignment of the generated audio samples.
@@ 5.2 Results
SpecVQGAN 26.63 5.57 3.30 0.1336 0.1422 0.2851 379M
Im2wav 16.87 5.94 2.53 0.4001 0.1310 0.2763 365M
Diff-Foley 21.96 6.46 3.15 0.4010 0.1571 0.2059 859M
Seeing&Hearing 20.72 6.58 2.34 0.5805 0.1668 0.1858 -
ReWaS (Ours) 15.24 2.16 2.78 0.4353 0.1149 0.3008 204M
Table 2: Performance comparison on VGGSound (Chen et al. 2020) with reproduced five seconds audio samples. “Energy” and “TP” denote energy MAE and number of the trainable parameters.
@@
CondFoleyGen 23.94 60.24 0.1520
ReWaS (Ours) 19.15 63.28 0.1398
Table 3: Performance comparison on Greatest Hits (Owens et al. 2016). We use material types as text prompts, while CondFoleyGen uses both reference audio and video as inputs.
Quantitative results. Table 2 shows the quantitative comparisons on the VGGSound. We note that category classes are used as text prompts in the VGGSound. We train 22M parameters for video projection to audio conditional control, and 182M parameters for fine-tuning the AudioLDM with our energy adapter. Since Seeing&Hearing is an optimization-based generation method, we did not report the training parameters. However, they consume twice the time for inference than ReWaS. Our ReWaS achieve the best performance on FD, FAD, energy MAE, and AV-align, showing competitive results in terms of MKL and CLAP scores. Especially, while we use only a quarter of training parameters compared to Diff-Foley, our method outperforms Diff-Foley on all metrics. CLAP scores illustrate the importance of text prompts for semantic alignment. Seeing&Hearing outperforms ReWaS in terms of MKL and CLAP score. However, we argue that Seeing&Hearing is heavily dependent on text prompt, since our method outperforms in terms of MAE and AV-align scores by a large margin. This achieved MAE score result by ReWaS also demonstrates the accuracy of our control prediction module, and generated outputs from ReWaS are most temporally closer to the real audio content.
In addition, we evaluate how the generated audio and the condition video are temporally aligned on Greatest Hits. The dataset distribution of Greatest Hits highly differs from the general audio samples; hence, we fine-tune ReWaS on the Greatest Hits training samples. Table 3 shows the results. ReWaS achieves the best AP and MAE, although ReWaS is not specially designed for Foley like CondFoleyGen.
@@
SpecVQGAN 2.76 2.50 2.64
Im2wav 2.97 3.18 3.01
Diff-Foley 2.89 2.97 2.98
@@
Table 4: Human evaluation of V2A methods on audio quality, audiovisual relevance, and temporal alignment with 5-scale MOS.
User study. The quantitative results are limited to measuring how the generated audio sounds realistic and aligned to the given video. To complement it, we conduct a human evaluation study to assess the subjective quality of the generated audio concerning the input video. We ask the human evaluators to evaluate the quality of the audio samples generated by SpecVQGAN, Im2wav, Diff-Foley, and ReWaS. Since Seeing&Hearing shows vulnerable performance in audio-visual alignment, we exclude it from the user study.
We evaluate three criteria: audio quality, relevance between audio and video, and temporal alignment. We use a five-point Likert scale to measure mean opinion score (MOS), where an ideal video with its ideal audio receives a rating of 5 across all criteria. We recruit human annotators via two separate channels: Amazon Mechanical Turk (AMT) and local hiring. We recruit 50 AMT annotators for each criterion, and each annotator evaluates five generated samples for each method (i.e., each annotator evaluates 20 audios). Locally hired 23 annotators evaluated 20 generated samples for each method and criterion. Surprisingly, ReWaS achieves the best in all categories with large margins as shown in Table 4. This subjective result is consistent with our quantitative and qualitative findings, further validating the effectiveness of ReWaS in generating high-quality, relevant, and temporally synchronized audio for the given video.
Figure 5: Qualitative comparison on VGGSound. Surprisingly, when the skateboarder jumps, only ReWaS succeded in detecting short transition (yellow box). Text prompt in is “skateboarding”.
Qualitative results. Figure 5 shows qualitative results in baselines and ReWaS. Given the skateboarding video, SpecVQGAN and Diff-Foley fail to generate the sound of skate wheels rolling on the floor. Although Im2wav generates that sound, it cannot capture a short transition. We also demonstrate the effectiveness of the text prompt in Figure 6 with CLAP similarity, when redundant frames exist. In this case, V2A methods also struggle to generate corresponding sound. However, ReWaS can effectively calibrate the semantics by user text prompt. Note that the prompt can also be longer and more general if desired by users. (e.g., “rally car swiftly navigates a turn on the racetrack.”)
@@
T & GT E from A AudioLDM-M 13.93 2.65 2.15 0.4497 0.1195
T & Est. E from V AudioLDM-M 15.24 2.16 2.78 0.4353 0.1149
T & Est. E from V Make-An-Audio 13.89 10.91 2.93 0.4237 0.1368
Table 5: Impact of the energy control’s quality on VGGSound. (1) Text and the ground-truth audio energy with AudioLDM backbone (upper bound), (2) Text and the estimated energy from the video with AudioLDM (our approach) and (3) with Make-An-Audio.
5.3 Discussion
The impact of the quality of the energy control. To verify the robustness of the energy prediction module, we compare the control by our video-to-energy prediction module and the energy directly extracted from the ground truth audio. Table 5 demonstrates that although we use the estimated energy, the quality of the generated audio is very similar to the audio samples controlled by the ground truth audio energy. (See the first row and second row of Table 5) It supports the idea that energy information is highly related to visual information, and is easy to estimate solely using video. We also compare the qualitative results of estimated energy controls with ground truth energy in the Section A.5.
T2A framework. We replace the AudioLDM-M backbone with Make-An-Audio (Huang et al. 2023b), which has fewer parameters than AudioLDM to validate the flexibility of our approach. Table 5 shows the results of the two backbones on VGGSound. Interestingly, ReWaS built upon Make-An-Audio achieves comparable performance to its AudioLDM-M counterpart, demonstrating the robustness of our framework (See the second row and third row of Table 5).
Figure 6: Effectiveness of text prompt. When videos contain both semantic and redundant frames, text prompts used in ReWaS calibrate the results. Text prompt in is “chicken clucking”
General text prompts. To examine the capability of ReWaS with more general text prompts, we generate audio samples with generative videos by KLING***https://kling.kuaishou.com/en. As shown in Figure 7, only our method can capture the visual information of the wave, namely, the sound getting louder as the wave crashes. We include more samples in the Section A.5 and demo.
Effectiveness of visual condition. In the Appendix, we show examples when energy control complementing temporal information. While AudioLDM suffers from inferior temporal alignment and limited sound generation that is mentioned in text prompts but not generated, ReWaS not only generates video-related sounds hidden in the text but also aligns the sound with the frames. This demonstrates the effectiveness of visual condition by ReWaS. More detailed discussion is in the Section A.5.
Figure 7: Audio generation with general text prompts.